# Graham’s rearrangement for dihedral groups

### v1

@@ 3 Structure Theorem
Following [4], we begin with $E^{\prime}=A$ and we assume $\dim(E^{\prime})>R$ (i.e. $s>0$). Then, as long as $\dim(E^{\prime})\geqslant R$, we could iteratively remove a dissociated subset $D^{\prime\prime\prime}$ of size $R/2$ and hence one, say $D^{\prime\prime}$, of size congruent to one modulo thirty-two, $R/4-31\leq|D^{\prime\prime}|\leq R/2$ such that $\pi_{2}(D^{\prime\prime})$ (i.e. the projection over $\mathbb{F}_{2}$) is constantly-zero or constantly-one. Also, if we divide $D^{\prime\prime}$ into two sets ${D^{o}}^{\prime\prime}=\{{d_{1}^{o}}^{\prime\prime},{d_{2}^{o}}^{\prime\prime}\dots,{d_{h+1}^{o}}^{\prime\prime}\}$ and ${D^{e}}^{\prime\prime}=\{{d_{1}^{e}}^{\prime\prime},{d_{2}^{e}}^{\prime\prime}\dots,{d_{h}^{e}}^{\prime\prime}\}$ then
$\delta^{\prime\prime}\vcentcolon={d_{1}^{o}}^{\prime\prime}\cdot{d_{1}^{e}}^{\prime\prime}\cdots{d_{h}^{o}}^{\prime\prime}\cdot{d_{h}^{e}}^{\prime\prime}\cdot{d_{h+1}^{e}}^{\prime\prime}$
does not depend on the orderings on ${D^{o}}^{\prime\prime}$ and ${D^{e}}^{\prime\prime}$. Note that, deleting one element from ${D^{(o)}}^{\prime\prime}$, we can always assume that $\delta^{\prime\prime}$ is different from a given element $\bar{x}\in Dih_{p}$. This means that, as long as $\dim(E^{\prime})\geqslant R$, we could iteratively remove a dissociated subset $D^{\prime}$ of size multiple of thirty-two, $R/4-31\leq|D^{\prime}|\leq R/2$ such that $\pi_{2}(D^{\prime})$ is constantly-zero or constantly-one and the product of its elements (in a given order) is different from a chosen element $\bar{x}\in Dih_{p}$. Also, in this way, we obtain that $E^{\prime}$ is not fully contained in $\{(x,0):x\in\mathbb{F}_{p}\}$ whenever $A$ is not.
@@
where each $D_{j}^{\prime}$ is a dissociated set of size multiple of thirty-two $R/4-31\leq|D_{j}^{\prime}|\leq R/2$ and, for any $j\in L_{1}^{\prime}$, $D^{\prime}_{j}\subseteq\{(x,1):x\in\mathbb{F}_{p}\}$ and, for any $j\in L_{0}^{\prime}$, $D^{\prime}_{j}\subseteq\{(x,0):x\in\mathbb{F}_{p}\}$.
Given $j\in L_{1}^{\prime}$, we split $D_{j}^{\prime}$ in ${D_{j}^{o}}^{\prime}$ and ${D_{j}^{e}}^{\prime}$ so that $|{D_{j}^{o}}^{\prime}|=|{D_{j}^{e}}^{\prime}|$. Then we set
$\delta=\left(\prod_{j\in L_{1}^{\prime}}\prod_{{d^{o}}^{\prime}\in{D_{j}^{o}}^{\prime},{d^{e}}^{\prime}\in{D_{j}^{e}}^{\prime}}{d^{o}}^{\prime}\cdot{d^{e}}^{\prime}\right)\left(\prod_{j\in L_{0}^{\prime}}\prod_{d^{\prime}\in D^{\prime}_{j}}d^{\prime}\right)$
(which does not depend on the orderings on ${D_{j}^{o}}^{\prime}$ and ${D_{j}^{e}}^{\prime}$). Also, due to the previous consideration, we can assume that $\delta^{\prime}\not=(0,0)$.
Then Lemma 3.3 (applied to $E^{\prime}\cup\{\delta^{\prime}\}$) provides an automorphism $\phi\in Aut(Dih_{p})$ such that
@@
where $E$ satisfies property $(i)$ and each $D_{j}$ is a dissociated set of size multiple of thirty-two which satisfies properties $(ii)$ (with the stronger lower-bound $R/4-31$) and $(iii)$.
Setting $D_{j}^{o}=\phi({D_{j}^{o}}^{\prime})$, $D_{j}^{e}=\phi({D_{j}^{e}}^{\prime})$, and $\delta=\phi(\delta^{\prime})$, we have that, since $|D_{j}|$ is even for any $j$, $\delta$ is of the form $(z_{0},0)$ where $z_{0}\not=0$. Also, due to the choice of $\phi$,
$\pi_{1}(\phi(E\cup\{\delta^{\prime}\}))=\pi_{1}(E\cup\{\delta\})\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)$
@@
Following again [4], we can modify the decomposition to satisfy (v) as follows. We split $D_{1}^{o}$ and $D_{1}^{e}$ into $2$ parts $D_{1}^{o,(1)},D_{1}^{o,(2)}$ and $D_{1}^{e,(1)},D_{1}^{e,(2)}$ each of size $|D_{1}|/4\geq R/16-7$. Lemma 3.4 ensures that either $D_{1}^{o,(1)}\cup D_{1}^{e,(1)}\cup\{\delta\}$ or $D_{1}^{o,(2)}\cup D_{1}^{e,(2)}\cup\{\delta\}$ is dissociated. Setting $D_{1}^{(1)}=D_{1}^{o,(1)}\cup D_{1}^{e,(1)}$, without loss of generality, we can assume that $D_{1}^{(1)}\cup\{\delta\}$ is dissociated.
Then further split $D_{1}^{o,(1)}$ and $D_{1}^{e,(1)}$ into $2$ parts $D_{1}^{o,(3)},D_{1}^{o,(4)}$ and $D_{1}^{e,(3)},D_{1}^{e,(4)}$ each of size $|D_{1}|/8\geq R/32-3$, and we replace the sequence of sets $D_{1},\ldots,D_{s^{\prime}}$ by the sequence
@@
where $D_{1}^{(3)}=D_{1}^{o,(3)}\cup D_{1}^{e,(3)}$ and $D_{1}^{(4)}=D_{1}^{o,(4)}\cup D_{1}^{e,(4)}.$ For notation convenience, we rename this sequence as
@@
Note that, in the Structure Theorem, if $s>0$, we always obtain a decomposition where $|E|\geq R/2$.
We conclude this section by stating a structural kind of property of dissociated sets that will be important for the probabilistic section of this paper.
@@
Consider a dissociated set $D=D^{o}\cup D^{e}\subseteq Dih_{p}$ where the $\mathbb{F}_{2}$ component of each element of $D$ is $1$ and where $|D^{o}|=\lceil|D|/2\rceil$, $|D^{e}|=\lfloor|D|/2\rfloor$. Then for any
@@
where $D_{i}\neq D_{j}$, $d_{k},d_{k}^{\prime}\in D^{o}$ if $k$ is odd and $d_{k},d_{k}^{\prime}\in D^{e}$ if $k$ is even, we have
@@
Assume for the sake of contradiction that there exist
where $D_{i}\neq D_{j}$, $d_{k},d_{k}^{\prime}\in D^{o}$ if $k$ is odd and $d_{k},d_{k}^{\prime}\in D^{e}$ if $k$ is even, and we have
$d_{1}d_{2}\cdots d_{h}=d_{1}^{\prime}d_{2}^{\prime}\cdots d_{\ell}^{\prime}.$
Set $d_{k}=(x_{k},1)$ and $d_{k}^{\prime}=(x_{k}^{\prime},1)$, in components we have that
$(x_{1}-x_{2}+x_{3}-\dots(-1)^{h+1}x_{h},h\pmod{2})=(x_{1}^{\prime}-x_{2}^{\prime}+x_{3}^{\prime}-\dots(-1)^{\ell+1}x_{\ell}^{\prime},\ell\pmod{2}).$ (2)
This is only possible if $h\equiv\ell\pmod{2}$ and, if by moving everything to the lhs, we have that
$(x_{1}-x_{2}+x_{3}-\dots(-1)^{h+1}x_{h}-(x_{1}^{\prime}-x_{2}^{\prime}+x_{3}^{\prime}-\dots(-1)^{\ell+1}x_{\ell}^{\prime}),0)=(0,0).$ (3)
Here we note that the first component of the lhs of Equation (3) contains the same number of signs $+$ and --. Also, if some element $d=(x,1)$ appears both in the lhs and in the rhs of Equation (2), then we must have $x=x_{k}=x_{b}^{\prime}$ where $k\equiv b\pmod{2}$. Hence $x$ appears in Equation (3) with different signs. It means that after we erase these elements we still have the same number of signs $+$ and --. Hence Equation (3) becomes
@@
Setting $d_{k}^{\prime\prime}=(x_{k}^{\prime\prime},1)$ this implies that
$d_{1}^{\prime\prime}d_{2}^{\prime\prime}\cdot d_{t}^{\prime\prime}=id$
which is a contradiction of the fact that $D$ is dissociated. ∎
4 Ordering the Set $E$
Now, the proof proceeds mainly in two steps. Due to the Structure Theorem we can assume, without loss of generality, that
@@ 4 Ordering the Set $E$
where the $D_{j}$s are dissociated and satisfy the hypothesis of Theorem 3.5 and we have that $\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)$. We recall that, given $d\in Dih_{p}$, we have set $\pi_{1}(d)$ and $\pi_{2}(d)$ the projections of $d$ over, respectively, $\mathbb{F}_{p}$ and $\mathbb{F}_{2}$. We begin by ordering the set $E$ and, for this purpose, we set some new notations. We define
$P=\left\{x\in E:\pi_{2}(x)=0\mbox{ and }\pi_{1}(x)\in\left(0,\frac{p}{90(|E|+1)}\right)\right\},$
$N=\left\{x\in E:\pi_{2}(x)=0\mbox{ and }\pi_{1}(x)\in\left(-\frac{p}{90(|E|+1)},0\right)\right\},$
$S=\{x\in E:\pi_{2}(x)=1\}$
where $S$ is not empty due to the Structure Theorem.
Also we split $S$ into $S=S_{e}\cup S_{o}$ so that $|S_{e}|=\lceil|S|/2\rceil$, $|S_{o}|=\lfloor|S|/2\rfloor$ and
$\forall x\in S_{e},\forall y\in S_{o},\ \pi_{1}(x)>\pi_{1}(y).$
Finally, given an ordering $\mathbf{x}=x_{0},x_{1},\dots,x_{m}$ we consider the set of partial products
@@
Now we show that, for any family of subsets $Y_{1},\ldots,Y_{m}\subseteq Dih_{p}$, we can order $E$ as follows:
@@
Let $p$ be large enough. Consider $\delta=(z_{0},0)\in Dih_{p}$ where $z_{0}>0$ and subsets $Y_{1},\ldots,Y_{m}\subseteq Dih_{p}$. Given a finite subset $E$ such that $\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)\subseteq Dih_{p}$, $E\not\subseteq\{(x,0):x\in\mathbb{F}_{p}\}$, consider $P,N,S_{e}$ and $S_{o}$ defined as above. Then there are orderings $\mathbf{p}$ of $P$, $\mathbf{n}$ of $N$, $\mathbf{s}_{e}=(s_{0},s_{2},\dots,s_{2\ell})$ of $S_{e}$ (here $\ell=\lceil|S|/2\rceil-1$) and $\mathbf{s}_{o}=(s_{1},s_{3},\dots,s_{2h-1})$ of $S_{o}$ (here $h=\lfloor|S|/2\rfloor$) such that, setting
$\mathbf{x}=\mathbf{p}\ s_{0}\ \mathbf{n}\ s_{1}s_{2}\dots,s_{\ell+h},$
$\delta,\mathbf{x}$ is a sequencing, and we have
$|\delta\cdot\operatorname{IS}(\mathbf{x})\cap Y_{j}|\leqslant 4\inf_{L\in\mathbb{N}}\left(\frac{|Y_{j}|}{L}+L+2\sum_{i=1}^{j-1}|Y_{i}|\right)$ (4)
for all $1\leqslant j\leqslant m$.
@@
First, note that any such ordering is a sequencing. Also, since
$\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)\subseteq Dih_{p},$
@@
Now, we will construct the sequences $\mathbf{x}=\mathbf{p}\ s_{0}\ \mathbf{n}\ s_{1}s_{2}\dots,s_{\ell+h}=x_{1}x_{2}\dots x_{|P|+|N|+|S|}$ from the larger indices to the smaller indices (with respect to $\mathbf{x}$). Here we set $x_{i}=(z_{i},\epsilon_{i})$ where $\epsilon_{i}\in\{0,1\}$ according to whether $x_{i}\in P\cup N$ or $x_{i}\in S$.
@@
    if $k>|P|+|N|+1$ and $k-(|P|+|N|+1)$ is even or if $k=|P|+1$, we need to choose an $x_{k}\in S_{e}$;
@@
$\nu_{k}:=z_{0}+\sum_{x\in P^{k}\cup S_{e}^{k}}\pi_{1}(x)-\sum_{x\in N^{k}\cup S_{o}^{k}}\pi_{1}(x).$
Note that, for any ordering $\mathbf{x}=x_{1}x_{2}\dots x_{|P|+|N|+|S|}$, $\nu_{k}=\pi_{1}\left(\delta\prod_{i=1}^{k}x_{i}\right)$.
@@
alternatively in $S_{e}$ and in $S_{o}$ (according to the parity of $k-(|P|+|N|+1)$) and we need to choose $x_{k}\in S_{o}$. Following the idea of Proposition 4.1 of [4], we choose $x_{k}$ as follows:
  1. 1.
    If there is some $x^{*}=(z^{*},1)\in S_{o}^{k}$ such that $(\nu_{k}+z^{*},1)\notin\cup_{j}Y_{j}$, then choose $x_{k}$ to be this $x^{*}$ and say that the current step is a skip-step for $S_{o}$.
  2. 2.
    Now, consider the case where $(\nu_{k}+\pi_{1}(S_{o}^{k}),1)\subseteq\cup_{j}Y_{j}$. Let $i$ be minimal such that $(\nu_{k}+\pi_{1}(S_{o}^{k}),1)$ intersects $Y_{i}$, and say that the current step is an $i$-step for $S_{o}$. If $(\nu_{k}+\pi_{1}(S_{o}^{k}),1)\subseteq Y_{i}$, then we set

### v2

@@ 3. Semidirect products
Following [4], we begin with $E^{\prime}=A$ and assume that $\dim(E^{\prime})>R$ (i.e. $s>0$). Then, as long as $\dim(E^{\prime})\geq R$, we can iteratively remove a dissociated subset $D^{\prime\prime\prime}$ of size $R/2$ and hence one, say $D^{\prime\prime}$, whose size is in the interval [R/431,R/2][R/4-31,R/2] and congruent to one modulo thirty-two, and such that $\pi_{2}(D^{\prime\prime})$ (i.e. the projection over $H$) is constantly in $H_{0}$ (i.e. the set $h\in H:\varphi(h)=id$) or constantly in $H_{1}$ (i.e. the set $h\in H:\varphi(h)=-id$). Also, if we divide $D^{\prime\prime}$ into two sets ${D^{o}}^{\prime\prime}=\{{d_{1}^{o}}^{\prime\prime},{d_{2}^{o}}^{\prime\prime}\dots,{d_{h+1}^{o}}^{\prime\prime}\}$ and ${D^{e}}^{\prime\prime}=\{{d_{1}^{e}}^{\prime\prime},{d_{2}^{e}}^{\prime\prime}\dots,{d_{h}^{e}}^{\prime\prime}\}$, then
$\delta^{\prime\prime}\vcentcolon={d_{1}^{o}}^{\prime\prime}\cdot{d_{1}^{e}}^{\prime\prime}\cdots{d_{h}^{o}}^{\prime\prime}\cdot{d_{h}^{e}}^{\prime\prime}\cdot{d_{h+1}^{o}}^{\prime\prime}$
does not depend on the orderings on $D^{o}$$^{\prime\prime}$ and $D^{e}$$^{\prime\prime}$. Note that up to deleting one element from $D^{o}$$^{\prime\prime}$, we can always assume that $\delta^{\prime\prime}$ is different from a given element $\bar{x}\in G$. This means that as long as $\dim(E^{\prime})\geq R$, we can iteratively remove a dissociated subset $D^{\prime}$, whose size is in the interval [R/432,R/2][R/4-32,R/2] and is a multiple of thirty-two, and such that $\pi_{2}(D^{\prime})$ is constantly in $H_{0}$ or constantly in $H_{1}$ and the product of its elements (in a given order) is different from a chosen element $\bar{x}\in G$.
@@
where each $D_{j}^{\prime}$ is a dissociated set whose size is in the interval $R/4-32\leq|D_{j}^{\prime}|\leq R/2$ and is a multiple of thirty-two, for any $j\in L_{1}^{\prime}$ we have that $D^{\prime}_{j}\subseteq\{(x,a):x\in\mathbb{Z}_{p},a\in H_{1}\}$, and for any $j\in L_{0}^{\prime}$ we have that $D^{\prime}_{j}\subseteq\{(x,a):x\in\mathbb{Z}_{p},a\in H_{0}\}$.
Given $j\in L_{1}^{\prime}$, we split $D_{j}^{\prime}$ into $D_{j}^{o}$$^{\prime}$ and $D_{j}^{e}$$^{\prime}$ so that $|{D_{j}^{o}}^{\prime}|=|{D_{j}^{e}}^{\prime}|$. Then we set
$\delta^{\prime}=\left(\prod_{j\in L_{1}^{\prime}}\prod_{{d^{o}}^{\prime}\in{D_{j}^{o}}^{\prime},{d^{e}}^{\prime}\in{D_{j}^{e}}^{\prime}}{d^{o}}^{\prime}\cdot{d^{e}}^{\prime}\right)\left(\prod_{j\in L_{0}^{\prime}}\prod_{d^{\prime}\in D^{\prime}_{j}}d^{\prime}\right),$
which does not depend on the orderings on $D_{j}^{o}$$^{\prime}$ and $D_{j}^{e}$$^{\prime}$. Also, due to the previous consideration, we can assume that $\delta^{\prime}\not=(0,0)$.
Then Lemma 3.3 (applied to $E^{\prime}\cup\{\delta^{\prime}\}$) provides an automorphism $\phi\in Aut(G)$ such that
@@
where $E$ satisfies property $(i)$ and each $D_{j}$ is a dissociated set of size a multiple of thirty-two which satisfies properties $(ii)$ (with the stronger lower bound R/432R/4-32) and $(iii)$.
Setting $D_{j}^{o}=\phi({D_{j}^{o}}^{\prime})$, $D_{j}^{e}=\phi({D_{j}^{e}}^{\prime})$, and $\delta=\phi(\delta^{\prime})$, we have that, since $|D_{j}|$ is even for any $j$, $\delta$ is of the form $(z_{0},a_{0})$ where $z_{0}\not=0$ and $a_{0}\in H_{0}$. Also, due to the choice of $\phi$,
$\pi_{1}(\phi(E^{\prime}\cup\{\delta^{\prime}\}))=\pi_{1}(E\cup\{\delta\})\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)$
@@
Following again [4], we can modify the decomposition to satisfy (v) as follows. We split $D_{1}^{o}$ and $D_{1}^{e}$ into $2$ parts $D_{1}^{o,(1)},D_{1}^{o,(2)}$ and $D_{1}^{e,(1)},D_{1}^{e,(2)}$ each of size $|D_{1}|/4\geq R/16-8$. Lemma 3.4 ensures that either $D_{1}^{o,(1)}\cup D_{1}^{e,(1)}\cup\{\delta\}$ or $D_{1}^{o,(2)}\cup D_{1}^{e,(2)}\cup\{\delta\}$ is dissociated. Setting $D_{1}^{(1)}=D_{1}^{o,(1)}\cup D_{1}^{e,(1)}$, without loss of generality, we can assume that $D_{1}^{(1)}\cup\{\delta\}$ is dissociated.
Then further split $D_{1}^{o,(1)}$ and $D_{1}^{e,(1)}$ into $2$ parts $D_{1}^{o,(3)},D_{1}^{o,(4)}$ and $D_{1}^{e,(3)},D_{1}^{e,(4)}$ each of size $|D_{1}|/8\geq R/32-4$. Replace the sequence of sets $D_{1},\ldots,D_{s^{\prime}}$ by the sequence
@@
where $D_{1}^{(2)}=D_{1}^{o,(2)}\cup D_{1}^{e,(2)}$, $D_{1}^{(3)}=D_{1}^{o,(3)}\cup D_{1}^{e,(3)}$, and $D_{1}^{(4)}=D_{1}^{o,(4)}\cup D_{1}^{e,(4)}.$ For notational convenience, we rename this sequence as
@@
Note that, in the Structure Theorem 3.5, if $s>0$ then we always obtain a decomposition where $|E|\geq R/2$.
We conclude this section by stating a structural property of dissociated sets that is important for the probabilistic section of this paper.
@@
Consider a dissociated set $D=D^{o}\cup D^{e}\subseteq G$ where $|D^{o}|=\lceil|D|/2\rceil$ and $|D^{e}|=\lfloor|D|/2\rfloor$ for $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian, $\varphi$ satisfies property $(\star)$, and the $H$ component of each element of $D$ is in $H_{1}$. Then for any
@@
where $D_{i}\neq D_{j}$, $d_{k},d_{k}^{\prime}\in D^{o}$ if $k$ is odd, and $d_{k},d_{k}^{\prime}\in D^{e}$ if $k$ is even, we have
@@
Assume for the sake of contradiction that there exist $D_{i}=\{d_{1},\dots,d_{h}\},\ D_{j}=\{d_{1}^{\prime},\dots,d_{\ell}^{\prime}\}\subseteq D$ where $D_{i}\neq D_{j}$, $d_{k},d_{k}^{\prime}\in D^{o}$ if $k$ is odd, $d_{k},d_{k}^{\prime}\in D^{e}$ if $k$ is even, and $d_{1}d_{2}\cdots d_{h}=d_{1}^{\prime}d_{2}^{\prime}\cdots d_{\ell}^{\prime}.$ Setting $d_{k}=(x_{k},a_{k})$ and $d_{k}^{\prime}=(x_{k}^{\prime},a_{k}^{\prime})$, in components we have that
(4) $(x_{1}-x_{2}+x_{3}-\dots(-1)^{h+1}x_{h},\sum_{i=1}^{h}a_{i})=(x_{1}^{\prime}-x_{2}^{\prime}+x_{3}^{\prime}-\dots(-1)^{\ell+1}x_{\ell}^{\prime},\sum_{i=1}^{\ell}a_{i}^{\prime}).$
This is only possible if $h\equiv\ell\pmod{2}$ since $(-id)^{h}=\varphi(\sum_{i=1}^{h}a_{i})=\varphi(\sum_{i=1}^{\ell}a^{\prime}_{i})=(-id)^{\ell}$ and by moving everything to the left-hand side, we get
(5) $(x_{1}-x_{2}+x_{3}-\dots(-1)^{h+1}x_{h}-(x_{1}^{\prime}-x_{2}^{\prime}+x_{3}^{\prime}-\dots(-1)^{\ell+1}x_{\ell}^{\prime}),0)=(0,0).$
Here we note that the first component of the left-hand side of Equation (5) contains the same number of ++ and - signs. Also, if some element $d=(x,a)$ appears both in both the left-hand and right-hand sides of Equation (4), then we must have $x=x_{k}=x_{b}^{\prime}$ where $k\equiv b\pmod{2}$. Hence $x$ appears in Equation (5) with different signs, so we still have the same number of ++ and - signs after erasing these repeated elements. Thus Equation (5) becomes
@@
Setting $d_{k}^{\prime\prime}=(x_{k}^{\prime\prime},\pm a_{k}^{\prime\prime})$ (according to whether $d_{k}^{\prime\prime}$ comes from the right- or left-hand side of Equation (4)) this implies that
$d_{1}^{\prime\prime}d_{2}^{\prime\prime}\cdots d_{t}^{\prime\prime}=id.$
Since $d_{k}^{\prime-1}=(x_{k}^{\prime},-a_{k}^{\prime})$, the previous relation is in contradiction with the fact that $D$ is dissociated. ∎
4. Ordering the Set $E$ and the Dissociated Sets
Now, the proof proceeds mainly in two steps. Due to the Structure Theorem 3.5, we can assume without loss of generality that
@@ 4. Ordering the Set $E$ and the Dissociated Sets
where the $D_{j}$s are dissociated and satisfy the hypotheses of Theorem 3.5 and we have that $\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)$. Recall that, given $d\in\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian and $\varphi$ satisfies property $(\star)$, we let $\pi_{1}(d)$ and $\pi_{2}(d)$ be the projections of $d$ over $\mathbb{Z}_{p}$ and $H$, respectively. We begin by ordering the set $E$ and, for this purpose, set some new notations. We split $E$ into four sets:
$P=\left\{x\in E:\varphi(\pi_{2}(x))=id\mbox{ and }\pi_{1}(x)\in\left(0,\frac{p}{90(|E|+1)}\right)\right\},$
(6) $N=\left\{x\in E:\varphi(\pi_{2}(x))=id\mbox{ and }\pi_{1}(x)\in\left(-\frac{p}{90(|E|+1)},0\right)\right\},$
$Z=\left\{x\in E:\varphi(\pi_{2}(x))=id\mbox{ and }\pi_{1}(x)=0\right\},$
$S=\{x\in E:\varphi(\pi_{2}(x))=-id\}.$
Also, if $S$ is nonempty, we split it into $S=S_{e}\cup S_{o}$ such that $|S_{e}|=\lceil|S|/2\rceil$, $|S_{o}|=\lfloor|S|/2\rfloor$, and
(7) $\forall x\in S_{e},\forall y\in S_{o},\ \pi_{1}(x)\geq\pi_{1}(y).$
Finally, given an ordering $\mathbf{x}=x_{0},x_{1},\dots,x_{m}$, we consider the set of partial products
@@
By the Structure Theorem 3.5, we have that $\delta=(z_{0},a_{0})$ where $\varphi(a_{0})=id$ and $z_{0}\neq 0$.
We consider the case where $S$ is empty in the remark below.
Remark 4.1.
According to Lemma 3.4, up to absorbing, in the proof of Theorem 3.5, one element of $Z$ into the dissociated sets $D_{j}$s, we can impose that $\sum_{z\in Z}z\neq id$.
Now we show that, for any family of subsets $Y_{1},\ldots,Y_{m}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$, we can order $E$ as long as $S\neq\emptyset$ as follows:
@@
Let $p$ be large enough. Consider $\delta=(z_{0},0)\in\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $z_{0}>0$ and subsets $Y_{1},\ldots,Y_{m}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$. Given a finite subset $E$ such that $\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$, consider $P,N,Z,S,S_{e}$ and $S_{o}$ defined as above where $S\neq\emptyset$. Then there are orderings $\mathbf{p}$ of $P$, $\mathbf{n}$ of $N$, $\mathbf{z}$ of $Z$, $\mathbf{s}_{e}=(s_{0},s_{2},\dots,s_{2\ell})$ of $S_{e}$ (here $\ell=\lceil|S|/2\rceil-1$) and $\mathbf{s}_{o}=(s_{1},s_{3},\dots,s_{2h-1})$ of $S_{o}$ (here $h=\lfloor|S|/2\rfloor$) such that, setting
(8) $\mathbf{x}=\mathbf{z}\ \mathbf{p}\ s_{0}\ \mathbf{n}\ s_{1}s_{2}\dots,s_{\ell+h},$
we have that $\delta,\mathbf{x}$ is a sequencing and
(9) $|\delta\cdot\operatorname{IS}(\mathbf{x})\cap Y_{j}|\leq 4\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)+|H|$
for all $1\leq j\leq m$.
@@
First, note that since $H$ is strongly sequenceable, there exists a sequencing $\mathbf{z}$ of $Z$. Then for such an ordering we can easily see that any ordering defined in (8) is a sequencing for $E$. Also, since
$\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H,$
@@
Now, we will construct the sequences $\mathbf{x}=\mathbf{z}\ \mathbf{p}\ s_{0}\ \mathbf{n}\ s_{1}s_{2}\dots,s_{\ell+h}=\mathbf{z}\ x_{1}x_{2}\dots x_{|P|+|N|+|S|}$ from the larger indices to the smaller indices (with respect to $\mathbf{x}$). We set $x_{i}=(z_{i},\epsilon_{i})$ where $\epsilon_{i}\in\{0,1\}$ according to whether $x_{i}\in P\cup N$ or $x_{i}\in S$, respectively.
@@
    if $k>|P|+|N|+1$ and k(|P|+|N|+1)k-(|P|+|N|+1) is even or if $k=|P|+1$, we need to choose an $x_{k}\in S_{e}$;
@@
$\nu_{k}:=\delta+\sum_{x\in P^{k}\cup S_{e}^{k}}x-\sum_{x\in N^{k}\cup S_{o}^{k}}x.$
Note that, for any ordering $\mathbf{x}=\mathbf{z}\ x_{1}x_{2}\dots x_{|P|+|N|+|S|}$, we have that $\nu_{k}=\delta\cdot\prod_{z\in Z}z\cdot\prod_{i=1}^{k}x_{i}$.
@@
alternatively in $S_{e}$ and in $S_{o}$ (according to the parity of k(|P|+|N|+1)k-(|P|+|N|+1)) and we need to choose $x_{k}\in S_{o}$. Following the idea of Proposition 4.1 of [4], we choose $x_{k}$ as follows:
  1. (1)
    If there is some $x^{*}\in S_{o}^{k}$ such that $\nu_{k}+x^{*}\notin\cup_{j}Y_{j}$, then choose $x_{k}$ to be this $x^{*}$ and say that the current step is a skip-step for $S_{o}$.
  2. (2)
    Now, consider the case where $\nu_{k}+S_{o}^{k}\subseteq\cup_{j}Y_{j}$. Let $i$ be minimal such that $\nu_{k}+S_{o}^{k}$ intersects $Y_{i}$, and say that the current step is an $i$-step for $S_{o}$. If $\nu_{k}+S_{o}^{k}\subseteq Y_{i}$, then we set